OpenStudy (emcheshhhh):

Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (wanell):

PART A) The slopes of f(x) and g(x) are the same. PART B) The equation for line f(x) y = 4x + b -1 = 4(0) + b -1 = b y = 4x - 1 The function G(x) has a bigger y-intercept. The y-intercept of g(x) is 3, while the y-intercept of f(x) is -1. does this help